Circular No. 6050

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS T. Belloni, Astronomical Institute ‘Anton Pannekoek’, University of

Amsterdam, writes: “We have analyzed the ROSAT archive data for the source 1ES 1649–403 (IAUC 6048), which has been proposed as a counterpart of the BATSE x-ray nova in Scorpius (IAUC 6046). The ROSAT PSPC observation was made over an 18-day period in 1991 Feb./Mar.; the source is clearly detected at a large off-axis angle, which, however, does not affect the results of our analysis. The source shows strong variability; at the beginning (Feb. 28) and at the end (Mar. 18) of the observations, it is at a constant level of ∼ 0.1 count/s (total 0.1- to 2.4-keV flux ∼ 10^{12} erg cm^{−2} s^{−1}), while in the middle of the observation its count rate is increasing over a period of 12 000 s, reaching an apparent peak of 0.9 count/s on Mar. 3. The x-ray spectrum shows a considerable number of counts below 0.4 keV, indicative of a negligible interstellar absorption. A fit to a thermal spectrum gives N_{H} < 3 × 10^{19} cm^{−2} and kT > 1 keV. The absence of a substantial absorption suggests that the source is located nearby, and the observed time variability indicates that it might be a flare star. We therefore conclude that this source cannot be a counterpart for the recently-discovered BATSE x-ray nova.”

C. Bailyn, S. Jogee, and J. Orosz, Department of Astronomy, Yale University, report: “CCD imaging of the field of GRO J1655–40 (IAUC 6046) with the Cerro Tololo 0.9-m telescope (+ 2048 Tek CCD) has revealed a candidate optical counterpart for the high-energy nova. The optical nova is located at α = 16^{h}54^{m}00^{s}, δ = −39^{o}50′44′′ (equinox 2000.0; ± 2′′); its magnitude and colors on Aug. 11.038 UT were V = 14.2, B − V = +1.1, V − R = +0.8, R − I = +0.7 (± 0.1 mag). The digitized U.K. Schmidt Sky Survey shows only an object of V ∼ 18 at this location. This object may be a near neighbor of the nova that is visible on the CCD images (and not associated with the quiescent nova). The red color of the nova suggests that, if it is associated with GRO J1655–40, it is significantly absorbed. The candidate also shows an excess of 0.45 mag in a 7.5-nm-wide filter centered on Hα, relative to other stars in the field. We have also imaged the field of 1ES 1649–403 (IAUC 6048) and find no sources significantly brighter than they appear on the sky survey. Cerro Tololo currently has no spectroscopic instruments available, so spectroscopic confirmation from other observatories is urgently requested. Finding charts may be obtained from orosz@astro.yale.edu.”

1994 August 11 Daniel W. E. Green